OpenStudy (anonymous):

what is cos (11π/3)? explain

OpenStudy (anonymous):

1/2 11pi/3-2pi= 5pi/3 =5x180/3= 300 cos 300 = 1/2

OpenStudy (anonymous):

thanks so much :) is that the answer 1/2?

OpenStudy (anonymous):

yes. you're very welcome

OpenStudy (anonymous):

oh ok just making sure. when i did cos 300 i got a different answer but then i realized it was in radians instead of degrees anyways thanks again! :)